# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 6 Discussion
This paper introduces a model-oriented framework for defining distances between graphs within a given class, where each graph is treated as a statistical model and the collection of graphs is organized as a poset based on model inclusion. The resulting metric, derived from the shortest path through neighboring models, provides a principled way to measure distance across various graph classes including causal DAGs, probabilistic CPDAGs, and causal MPDAGs.
Our work opens several avenues for future research. First, for efficient computation of the model-oriented distance for MPDAGs, it is important to develop an explicit graphical characterization of the neighborhood structure. Second, one can consider a weighted version of the shortest path, where a weight is assigned to each covering pair in the poset. One possible approach for selecting these weights is to base them on information-theoretic notions, as was done in the context of distances between two partitions (Meila, 2005). Moreover, computational efficiency in executing the A* algorithm can be substantially improved through better heuristics, parallelization, and memory management. The authors are currently pursuing a separate project to reduce the time it takes to compute the model-oriented distance. Finally, the proposed framework is applicable to other graph classes including local independence graphs for capturing dependence structure of events that occur in time (Didelez, 2008; Mogensen and Hansen, 2020), maximal ancestral graphs (Richardson and Spirtes, 2002) and partial ancestral graph (Spirtes, Glymour and Scheines, 2000) that account for model dependencies in the presence of latent variables.
@@
[Acknowledgments] The authors thank undergraduate students Arnav Mazumder, Jian Kang, and Carter Lembo at the University of Washington for helping with the implementations of our algorithms.
@@ A.1 Injective model maps
We show that the models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$, as defined in Section˜2.2, are injective in the sense that no two different graphs represent the same model.
Assumption 1.
@@
Theorem 5.
Under ˜1, we have the following results.
@@
    Further, under ˜1, there exists $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ that is faithful to $\mathcal{G}$, i.e.,
@@
    Under ˜1, there also exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D})$ such that
@@ A.2 Graphical characterization of model containment
Proof of Proposition˜6.
By definition of the causal model associated with a DAG (see Section˜2.2.2), if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction by contradiction. Suppose $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ but $\mathcal{G}_{1}$ contains an edge $i\rightarrow j$ that is absent from $\mathcal{G}_{2}$. Recall that $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$ is the sample space. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. We can choose a law $P\in\mathcal{M}(\mathcal{G}_{1})$ of $X(\cdot)$ such that $\operatorname{\mathbb{E}}[X_{j}(x)]=x_{i}$ holds for every $x\in\mathcal{X}$. However, under $\mathcal{M}(\mathcal{G}_{2})$, $\operatorname{\mathbb{E}}[X_{j}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{G}_{2}}(j)}$ with $i\notin\operatorname{Pa}_{\mathcal{G}_{2}}(j)$, which is a contradiction. ∎
Proof of Proposition˜9.
Recall from Section˜2.2 that $\mathcal{M}(\mathcal{G}_{i}):=\cup_{\mathcal{D}\in\mathcal{G}_{i}}\mathcal{M}(\mathcal{D})$ for $i=1,2$, where $\mathcal{D}\in\mathcal{G}_{i}$ is any DAG represented by the MPDAG $\mathcal{G}_{i}$. It follows from definition of $\mathcal{M}(\mathcal{D})$ and Proposition˜6 that, if for every $\mathcal{D}_{1}\in\mathcal{G}_{1}$, there exists a certain $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. Fix any $\mathcal{D}_{1}\in\mathcal{G}_{1}$. We want to show that $\mathcal{D}_{1}$ is a subgraph of some $\mathcal{D}\in\mathcal{G}_{2}$. We can choose a law $P\in\mathcal{M}(\mathcal{D}_{1})$ of $X(\cdot)$ such that for every $i\in V$, $x\in\mathcal{X}$, it holds that
@@
where $\beta_{i}\in\mathbb{R}^{|\operatorname{Pa}_{\mathcal{D}_{1}}(i)|}$ is a fixed vector with non-zero coordinates. Since $P\in\mathcal{M}(\mathcal{D}_{1})\subseteq\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})=\cup_{\mathcal{D}\in\mathcal{G}_{2}}\mathcal{M}(\mathcal{D})$, there must exists $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $P\in\mathcal{M}(\mathcal{D}_{2})$. By definition, under $\mathcal{M}(\mathcal{D}_{2})$, $\operatorname{\mathbb{E}}[X_{i}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{D}_{2}}(i)}$. Hence, Eq.˜5 implies that $\operatorname{Pa}_{\mathcal{D}_{1}}(i)\subseteq\operatorname{Pa}_{\mathcal{D}_{2}}(i)$ for every $i\in V$. That is, $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$. ∎
@@ B.1 CPDAG
    Each connected component induced by undirected edges is chordal;
@@ B.2 MPDAG
    For each connected component $\mathrm{C}$ induced by undirected edges, the skeleton of $\mathcal{G}_{\mathrm{C}}$ is chordal;
@@ C.1 Simplification due to gradedness: proof of Proposition 1
Consider any path $p\in L(\mathcal{G}_{s},\mathcal{G}_{t})$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ where $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$. Let $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ be the sequence of graphs specifying this path. By definition, for each $i=2,\dots,k$, either $x_{i}$ covers $x_{i-1}$ or $x_{i-1}$ covers $x_{i}$. Since the poset is graded, we know that for each step, the rank increases or decreases by one, i.e. $|\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})|=1$. Note, by telescoping sum, we have:
$\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=\sum_{i=2}^{k}\mathrm{rank}(x_{i})-\mathrm{rank}(x_{i-1})$
@@
We have thus shown that the length of any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ is greater than or equal $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})$. Note that the lower-bound can be achieved by considering an upward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with the sequence $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ where $x_{i}$ covers $x_{i-1}$ for every $i=2,3,\dots,k$. Here, the inequality above becomes an equality since # $\text{ downward moves}=0$.
@@ Appendix D Proof of Theorem 2
Appendix E Properties of posets underlying different graph classes
@@ E.1 DAGs
    They directly follow from Proposition˜6 and the fact that given an arbitrary $\mathcal{G}\in\mathfrak{G}$, every subgraph of $\mathcal{G}$ is also in $\mathfrak{G}$.
@@
It is worth mentioning that when $|\mathrm{V}|>1$, $\mathcal{L}$ is neither a join semi-lattice nor upper semimodular. To see this, consider two DAGs that have an edge in opposite directions. They have no common upper bound by Proposition˜6.
@@ E.2 CPDAGs
By 2. of Proposition˜8, we have $\mathcal{G}_{1},\mathcal{G}_{2}\lessdot\mathcal{G}_{3},\mathcal{G}_{4}$. To see that $\mathcal{L}$ is not a join semi-lattice, observe that the upper bounds on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of $\mathcal{G}_{3},\mathcal{G}_{4}$ and the full graph $\mathcal{G}_{\hat{1}}$, where $\mathcal{G}_{3},\mathcal{G}_{4}\prec\mathcal{G}_{\hat{1}}$ but $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ are incomparable. Similarly, to see that $\mathcal{L}$ is not a meet semi-lattice, observe that the lower bounds on $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ consists of $\mathcal{G}_{1}$, $\mathcal{G}_{2}$ and the empty graph $\mathcal{G}_{\hat{0}}$. However, while $\mathcal{G}_{1},\mathcal{G}_{2}\succ\mathcal{G}_{\hat{0}}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are incomparable.
@@
By 2. of Proposition˜8, we know $\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}\gtrdot\mathcal{G}_{3}^{\prime}$. If $\mathcal{L}^{\prime}$ is upper semimodular, then there must exist a graph that covers both $\mathcal{G}_{1}^{\prime}$ and $\mathcal{G}_{2}^{\prime}$. Suppose $\mathcal{G}^{\prime}_{4}\gtrdot\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}$. By the same proposition, the skeleton of $\mathcal{G}^{\prime}_{4}$ must be either that of $\mathcal{G}_{5}^{\prime}$ or that of $\mathcal{G}_{6}^{\prime}$:
@@
Now we argue that $\mathcal{L}^{\prime}$ is not lower semimodular. For this purpose, consider again $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$ above. Proposition˜8 implies that the full graph covers $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. If $\mathcal{L}$ is lower semimodular, then there must exist $\mathcal{G}_{7}^{\prime}$ that is covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. By the same result, the skeleton of $\mathcal{G}_{7}^{\prime}$ is a 4-cycle and hence $\mathcal{G}_{7}^{\prime}$ must be either
@@
E.3 MPDAGs

### v2

@@ 6 Discussion
This paper introduces a model-oriented framework for defining a distance between graphs relative to a class of graphs in consideration. By treating each graph as a statistical (i.e., probabilistic or causal) model, we organize the class of graphs as a poset ordered by model inclusion. The poset induces a notion of neighborhood, which then yields a shortest path distance through neighbors as a metric in the space of graphs. By studying and exploiting the poset structures that are summarized in Table˜2, we develop A* algorithms for computing the model-oriented distance that can scale up to a moderate number of vertices.
By bridging graphical modeling with order theory, we open several avenues for future research. First, to compute the model-oriented distance between two MPDAGs, it remains an open problem to obtain an explicit graphical characterization of the neighborhood. Second, our distance can be extended to incorporate weights on the covering relations in the poset. In particular, one may introduce information-theoretic weights similar to those considered by Meila (2005) for partitions. Third, it is of interest to consider model selection in the spirit of Taeb, Bühlmann and Chandrasekaran (2024). Moreover, as an ongoing project, we are exploring ways to scale up the A* algorithm to larger graphs through better heuristic, upper bound, memory management and parallelization. Finally, our distance is in principle applicable to many other statistical graphs, such as maximal ancestral graphs (Richardson and Spirtes, 2002), partial ancestral graphs (Spirtes, Glymour and Scheines, 2000) and local independence graphs for event data (Didelez, 2008; Mogensen and Hansen, 2020).
@@
[Acknowledgments] The authors thank undergraduate students Jian Kang, Carter Lembo and Arnav Mazumder at the University of Washington for helping with the implementation of our algorithms.
@@ A.1 Injective model maps
We show that the models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$, as defined in Section˜2.3, are injective in the sense that no two different graphs represent the same model.
Assumption S1.
@@
Theorem S9.
Under ˜S1, we have the following results.
@@
    Further, under ˜S1, there exists $P\in\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}(\mathcal{G})$ that is faithful to $\mathcal{G}$, i.e.,
@@
    Under ˜S1, there also exists $P^{\prime}\in\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}(\mathcal{D})$ such that
@@ A.2 Graphical characterization of model containment
Proof of Proposition˜5.
By definition of the causal model associated with a DAG (see Section˜2.3.2), if $\mathcal{G}_{1}$ is a subgraph of $\mathcal{G}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction by contradiction. Suppose $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$ but $\mathcal{G}_{1}$ contains an edge $i\rightarrow j$ that is absent from $\mathcal{G}_{2}$. Recall that $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$ is the sample space. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. We can choose a law $P\in\mathcal{M}(\mathcal{G}_{1})$ of $X(\cdot)$ such that $\operatorname{\mathbb{E}}[X_{j}(x)]=x_{i}$ holds for every $x\in\mathcal{X}$. However, under $\mathcal{M}(\mathcal{G}_{2})$, $\operatorname{\mathbb{E}}[X_{j}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{G}_{2}}(j)}$ with $i\notin\operatorname{Pa}_{\mathcal{G}_{2}}(j)$, which is a contradiction. ∎
Proof of Proposition˜7.
Recall from Section˜2.3 that $\mathcal{M}(\mathcal{G}_{i}):=\cup_{\mathcal{D}\in\mathcal{G}_{i}}\mathcal{M}(\mathcal{D})$ for $i=1,2$, where $\mathcal{D}\in\mathcal{G}_{i}$ is any DAG represented by the MPDAG $\mathcal{G}_{i}$. It follows from definition of $\mathcal{M}(\mathcal{D})$ and Proposition˜5 that, if for every $\mathcal{D}_{1}\in\mathcal{G}_{1}$, there exists a certain $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$, then $\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})$. We now prove the other direction. Let us choose $\mathcal{X}_{i}=\mathbb{R}$ for every $i\in V$. Fix any $\mathcal{D}_{1}\in\mathcal{G}_{1}$. We want to show that $\mathcal{D}_{1}$ is a subgraph of some $\mathcal{D}\in\mathcal{G}_{2}$. We can choose a law $P\in\mathcal{M}(\mathcal{D}_{1})$ of $X(\cdot)$ such that for every $i\in V$, $x\in\mathcal{X}$, it holds that
@@
where $\beta_{i}\in\mathbb{R}^{|\operatorname{Pa}_{\mathcal{D}_{1}}(i)|}$ is a fixed vector with non-zero coordinates. Since $P\in\mathcal{M}(\mathcal{D}_{1})\subseteq\mathcal{M}(\mathcal{G}_{1})\subseteq\mathcal{M}(\mathcal{G}_{2})=\cup_{\mathcal{D}\in\mathcal{G}_{2}}\mathcal{M}(\mathcal{D})$, there must exist $\mathcal{D}_{2}\in\mathcal{G}_{2}$ such that $P\in\mathcal{M}(\mathcal{D}_{2})$. By definition, under $\mathcal{M}(\mathcal{D}_{2})$, $\operatorname{\mathbb{E}}[X_{i}(x)]$ (when it exists) can only depend on $x_{\operatorname{Pa}_{\mathcal{D}_{2}}(i)}$. Hence, Eq.˜8 implies that $\operatorname{Pa}_{\mathcal{D}_{1}}(i)\subseteq\operatorname{Pa}_{\mathcal{D}_{2}}(i)$ for every $i\in V$. That is, $\mathcal{D}_{1}$ is a subgraph of $\mathcal{D}_{2}$. ∎
@@ Appendix B Structural constraints of graph types
Recall that a partially directed acyclic graph (PDAG) is a simple graph with two types of edges, directed and undirected, such that there is no directed cycle.
@@ B.1 CPDAG
    Each connected component spanned by undirected edges is chordal;
@@ B.2 MPDAG
    For each connected component $\mathrm{C}$ spanned by undirected edges, the skeleton of $\mathcal{G}_{\mathrm{C}}$ ($G_{\mathrm{C}}$ is the subgraph induced by $\mathrm{C}$) is chordal;
@@ C.1 Simplification due to gradedness: proof of Proposition 1
Consider any path $p\in L(\mathcal{G}_{s},\mathcal{G}_{t})$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ where $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$. Let $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ be the sequence of graphs specifying this path. By definition, for each $i=2,\dots,k$, either $x_{i}$ covers $x_{i-1}$ or $x_{i-1}$ covers $x_{i}$. Since the poset is graded, we know that for each step, the rank increases or decreases by one, i.e., $|\operatorname{\mathrm{rank}}(x_{i})-\operatorname{\mathrm{rank}}(x_{i-1})|=1$. Note, by telescoping sum, we have:
$\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})=\sum_{i=2}^{k}\operatorname{\mathrm{rank}}(x_{i})-\operatorname{\mathrm{rank}}(x_{i-1})$
@@
We have thus shown that the length of any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ is greater than or equal $\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}_{s})$. Note that the lower-bound can be achieved by considering an upward path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ with the sequence $(x_{1}:=\mathcal{G}_{s},x_{2},\dots,x_{k}:=\mathcal{G}_{t})$ where $x_{i}$ covers $x_{i-1}$ for every $i=2,3,\dots,k$. Here, the inequality above becomes an equality since # $\text{ downward moves}=0$.
@@ Appendix D Proof of Theorem 2
Appendix E Posets properties for different graph classes
@@ E.1 DAGs
    They directly follow from Proposition˜5 and the fact that given an arbitrary $\mathcal{G}\in\mathfrak{G}$, every subgraph of $\mathcal{G}$ is also in $\mathfrak{G}$.
@@
It is worth mentioning that when $|\mathrm{V}|>1$, $\mathcal{L}$ is neither a join semi-lattice nor upper semimodular. To see this, consider two DAGs that have an edge in opposite directions. They have no common upper bound by Proposition˜5.
@@ E.2 CPDAGs
By 2. of Theorem˜5, we have $\mathcal{G}_{1},\mathcal{G}_{2}\lessdot\mathcal{G}_{3},\mathcal{G}_{4}$. To see that $\mathcal{L}$ is not a join semi-lattice, observe that the upper bounds on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ consists of $\mathcal{G}_{3},\mathcal{G}_{4}$ and the full graph $\mathcal{G}_{\hat{1}}$, where $\mathcal{G}_{3},\mathcal{G}_{4}\prec\mathcal{G}_{\hat{1}}$ but $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ are incomparable. Similarly, to see that $\mathcal{L}$ is not a meet semi-lattice, observe that the lower bounds on $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$ consists of $\mathcal{G}_{1}$, $\mathcal{G}_{2}$ and the empty graph $\mathcal{G}_{\hat{0}}$. However, while $\mathcal{G}_{1},\mathcal{G}_{2}\succ\mathcal{G}_{\hat{0}}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are incomparable.
@@
By 2. of Theorem˜5, we know $\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}\gtrdot\mathcal{G}_{3}^{\prime}$. If $\mathcal{L}^{\prime}$ is upper semimodular, then there must exist a graph that covers both $\mathcal{G}_{1}^{\prime}$ and $\mathcal{G}_{2}^{\prime}$. Suppose $\mathcal{G}^{\prime}_{4}\gtrdot\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime}$. By the same proposition, the skeleton of $\mathcal{G}^{\prime}_{4}$ must be either that of $\mathcal{G}_{5}^{\prime}$ or that of $\mathcal{G}_{6}^{\prime}$:
@@
Now we argue that $\mathcal{L}^{\prime}$ is not lower semimodular. For this purpose, consider again $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$ above. Theorem˜5 implies that the full graph covers $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. If $\mathcal{L}$ is lower semimodular, then there must exist $\mathcal{G}_{7}^{\prime}$ that is covered by both $\mathcal{G}_{5}^{\prime}$ and $\mathcal{G}_{6}^{\prime}$. By the same result, the skeleton of $\mathcal{G}_{7}^{\prime}$ is a 4-cycle and hence $\mathcal{G}_{7}^{\prime}$ must be either
@@
E.3 Polytree CPDAGs
Proof of Theorem˜6.
  1. 1.
    This directly follows from definition.
  2. 2.
[…]
    Using the facts above, we conclude the desired result. Let $\mathrm{F}=\{(i,j)\}$ be collection of vertices that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. By the analysis above, any v-structure in $\mathcal{G}_{1}$ is also present in $\mathcal{G}_{2}$. Furthermore, every v-structure $i\rightarrow k\leftarrow j$ in $\mathcal{G}_{2}$ is either present in $\mathcal{G}_{1}$ or at least one of the edges between pair of vertices $(i,k)$ and $(j,k)$ is missing. Consider any DAG $\mathcal{D}_{2}$ represented by $\mathcal{G}_{2}$. Let $\mathcal{D}_{1}$ be the DAG that is obtained by deleting the edges in $\mathcal{D}_{2}$ among pairs of vertices $\mathrm{F}$. By construction, $\mathcal{D}_{1}$ has the same skeleton as any DAG in $\mathcal{G}_{1}$. Since $\mathcal{D}_{1}$ is a polytree, every structure present in $\mathcal{D}_{1}$ or in $\mathcal{G}_{1}$ are also in $\mathcal{G}_{2}$. Suppose as a point of contradiction that there exists a v-structure $i\rightarrow k\leftarrow j$ that is in $\mathcal{D}_{1}$ and not in $\mathcal{G}_{1}$. Since this v-structure must also be present in $\mathcal{G}_{2}$, and from the previous statement, then one of the edges between pairs of vertices $(i,k)$ or $(j,k)$ must be missing in $\mathcal{G}_{1}$, which contradicts $\mathcal{D}_{1}$ and $\mathcal{G}_{1}$ have the same skeleton. A similar argument allows us to conclude that it is not possible that $\mathcal{G}_{1}$ has a v-structure that is not present in $\mathcal{D}_{1}$. Thus, $\mathcal{D}_{1}$ and $\mathcal{G}_{1}$ have the same skeleton and v-structures. We conclude that $\mathcal{D}_{1}$ must be a DAG represented by $\mathcal{G}_{1}$.
  3. 3.
    Let $\operatorname{\mathrm{rank}}(\cdot)$ be the number of edges. It suffices to show that $\operatorname{\mathrm{rank}}(\mathcal{G})+1=\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})$ holds for every $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ in $\mathcal{L}_{\mathrm{poly}}$. By the previous result, $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ implies $\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})\geq\operatorname{\mathrm{rank}}(\mathcal{G})+1$. For contradiction, suppose there exists $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ with $\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{rank}}(\mathcal{G}_{2})$. By the previous result, for every DAG $\mathcal{D}_{2}=(\mathrm{V},\mathrm{E}_{2})$ represented by $\mathcal{G}_{2}$, there exists a DAG $\mathcal{D}_{1}=(\mathrm{V},\mathrm{E}_{1})$ represented by $\mathcal{G}_{1}$ such that $\mathrm{E}_{1}\subseteq\mathrm{E}_{2}$. Since $\operatorname{\mathrm{rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{rank}}(\mathcal{G}_{2})$, there exists $\mathrm{E}_{3}$, satisfying $\mathrm{E}_{1}\subset\mathrm{E}_{3}\subset\mathrm{E}_{2}$. Let $\mathcal{G}_{3}$ be the CPDAG obtained by completing the DAG $\mathcal{D}_{3}=(\mathrm{V},\mathrm{E}_{3})$. By construction, we have $\mathcal{G}_{1}\preceq\mathcal{G}_{3}\preceq\mathcal{G}_{2}$ with $\mathcal{G}_{3}$ being distinct from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, contradicting $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
  4. 4.
    Consider any $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\in\mathfrak{G}_{\mathrm{poly}}$ with $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$ and $\mathcal{G}_{2}\lessdot\mathcal{G}_{3}$. We want to show that there is a graph that is covered by both $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Consider any DAG $\mathcal{D}_{3}=(\mathrm{V},\mathrm{E}_{3})\in[\mathcal{G}_{3}]$. By Result 2., there exist DAGs $\mathcal{D}_{1}=(\mathrm{V},\mathrm{E}_{1})\in[\mathcal{G}_{1}]$ and $\mathcal{D}_{2}=(\mathrm{V},\mathrm{E}_{2})\in[\mathcal{G}_{2}]$ with $\mathrm{E}_{1}\subseteq\mathrm{E}_{3}$ and $\mathrm{E}_{2}\subseteq\mathrm{E}_{3}$. Applying Result 3., we have $\mathrm{E}_{3}\setminus\mathrm{E}_{1}=e_{1}$ and $\mathrm{E}_{3}\setminus\mathrm{E}_{2}=e_{2}$ for two separate edges $e_{1},e_{2}$. Consider DAG $\tilde{\mathcal{D}}:=(\mathrm{V},\mathrm{E}_{1}\cap\mathrm{E}_{2})$ and let $\tilde{\mathcal{G}}$ be the CPDAG such that $\tilde{\mathcal{D}}\in[\tilde{\mathcal{G}}$]. By construction, $\tilde{\mathcal{G}}$ is a polytree CPDAG that satisfies $\tilde{\mathcal{G}}\preceq\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$. Since $\operatorname{\mathrm{rank}}(\tilde{\mathcal{G}})=\operatorname{\mathrm{rank}}(\mathcal{G}_{1})-1=\operatorname{\mathrm{rank}}(\mathcal{G}_{2})-1$, we have $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{2}$.
  5. 5.
    This follows from the previous result and Theorem˜2.
E.4 MPDAGs